OpenStudy (nanii215):

Given that F(x) = x^2 + 2, evaluate F(1) + F(5). A. 38 B. 30 C. 15

OpenStudy (jango_in_dtown):

plug in the value to get F(1)=1^2+2=3 and F(5)=5^2+2=27

OpenStudy (jango_in_dtown):

so F(1)+F(5)=3+27 = ???

OpenStudy (nanii215):

B 30?

OpenStudy (jango_in_dtown):

correct

OpenStudy (nanii215):

can you help me with more?

OpenStudy (jango_in_dtown):

Similar type of questionS?

OpenStudy (nanii215):

kinda.. i'll show you

OpenStudy (jango_in_dtown):

okay/

OpenStudy (nanii215):

If ƒ(x) = 2x^2 + 3, then which of the following represent ƒ(-1)? A. - 1 B. 1 C. 4 D. 5

OpenStudy (jango_in_dtown):

just plug in -1 in place of x f(-1)=2(-1)^2+3=2+3=?

OpenStudy (jango_in_dtown):

I think there is nothing much in the topic. just plug in the given x value and you can do it yourself

OpenStudy (nanii215):

Is it D. 5?

OpenStudy (jango_in_dtown):

yeah